OpenStudy (anonymous):

Identify the local maximum value for the function f (x) = x^6 - 6x^4 + 5, using the first derivative test.

OpenStudy (anonymous):

x = 2,-2

OpenStudy (anonymous):

also x=0

OpenStudy (anonymous):

Is this some kind of a joke?

OpenStudy (anonymous):

Why so???

OpenStudy (anonymous):

You tell me meanie :p You're giving me wrong answers on purpose -_-

OpenStudy (anonymous):

Evalute f'(x) = 0 and equate the equation. f(x) =x^6 - 6x^4 + 5 f'(x)=6x^5-24x^3 = 0 find the value of x .This will give you local maximum value

OpenStudy (anonymous):

i.e 6x^3(x^2 - 4) =0 find the possible value of x?

OpenStudy (anonymous):

the value of x from the equation will be x = 0,-2,+2. Am I correct?Otherwise you please explain it?

OpenStudy (anonymous):

Lol. The possible answers for this multiple choice question are 2, -2,4, or 0. Thats the reason I said that :/ Dont be mad at me :(